how many particles are in 2.66 of cu atoms?


1
Expert's answer
2020-10-12T13:35:26-0400

 mass of cu atoms = 2.66 g .


moles of cu atoms = 2.66 / 63.5 = 0.04188 mol .


number of atoms = moles of cu atom * 6.02 * 10^23 = 0.0418 * 6.02 * 1023 .


= 0.2516 * 10^23 .


=


2.5 * 10^22 atoms .